OpenStudy (anonymous):

6(7 -2p)=3(5p +1) HELP ASAP

OpenStudy (johnweldon1993):

Just distribute and simplify \[\large 6(7 - 2p)=3(5p + 1)\] Distribute the 6 and the 3 \[\large 42 - 12p = 15p + 3\] combine like terms \[\large 27p = 39\] Solve for p \[\large p = \frac{39}{27} = ?\]

OpenStudy (johnweldon1993):

But I do like the answer 42 lol seems sound :P